OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch side. A = sq. in.

OpenStudy (anonymous):

\[A = (s ^{2}\sqrt{3})/4\]

OpenStudy (anonymous):

\[A = (6^{2}\sqrt{3})/4\]

OpenStudy (anonymous):

\[A = (36\sqrt{3)})/4\]

OpenStudy (anonymous):

\[A = 9\sqrt{3}\]

OpenStudy (anonymous):

\[9\sqrt{3} \approx 15.6\]

OpenStudy (anonymous):

Use this formula whenever you need to find the area of an equilateral triangle.